# RULSurv: A probabilistic survival-based method for early censoring-aware prediction of remaining useful life in ball bearings

### v2

@@ 4.4 Data preprocessing
The XJTU-SY dataset is a tabular dataset, where the rows represent temporal feature values and the columns represent feature names. There are five bearings per operating condition and three operating conditions in total (see Table 2). As a preliminary preprocessing step, we convert the temporal dataset to a supervised learning dataset by computing a rolling average with adjustable window length ($w$) and lag ($l$) based on the operating condition. For bearings under high load (C1), we use $w=2$ and $l=-1$. For medium load (C2), we use $w=4$ and $l=-3$. For the light load (C3), we use $w=6$ and $l=-5$. This reflects the length of the event horizon and means that more samples are gathered from bearings running for longer, but also that the averaging window is larger and the lag is higher. The moving window averages both the features and the event times. This produces three supervised learning datasets, $\mathcal{D}_{C1}$ ($N=635$, $d=12$), $\mathcal{D}_{C2}$ ($N=1586$, $d=12$) and $\mathcal{D}_{C3}$ ($N=4930$, $d=12$), where $N$ is the number of samples and $d$ is the number of features.
@@ 4.5 Data censoring
Censored observations, i.e., where the event of failure is not observed, are common in an industrial setting. Some bearings may not have failed yet at the time of observation or they may have been decommissioned before failure occurred. Because the XJTU-SY dataset only includes complete observations, i.e., the failure time is always known, we introduce random and independent censoring by a fixed percentage, $C$, to simulate censoring in an industrial setting. The amount of censoring in an industrial setting can vary and depends on several factors, e.g., some bearings may be decommissioned in advance to avoid failure or to perform preventive maintenance. To evaluate our method under various, realistic amounts of censoring, we selected 25%, 50% and 75%, resulting in three censored datasets with the respective level of censoring. For each of these datasets, we randomly sampled the defined percentages of samples from the dataset $\mathcal{D}$, and designated those as censored. For example, to censor an observation, $i$, we calculate the time of censoring, $c_{i}$, as a random number between one and the observed event time, $e_{i}$, and use $c_{i}$ instead of $e_{i}$ as the time to event for observation $i$, i.e.$t_{i}=c_{i}$ and $\delta_{i}=0$. This indicates that the bearing was censored at some time between when it was recorded and when it actually failed. See B for a pseudocode implementation of the algorithm that automatically performs these changes.
Figure 4 shows the KM estimate of the survival function after random censoring has been applied to the individual datasets. The shaded area around the curve represent a empirical 95% confidence interval with upper and lower bounds. These bounds are the best/worst cases on the survival function given the observed data without any parametric assumptions. We calculate the bounds using the Greenwood formula [51], and they represent the predictive uncertainty and thus the consequences of censoring when making estimates. For the upper bound on the survival function, we assume that all censored bearings do not ever experience the event. For the lower bound, we assume that all censored bearings immediately experience the event after they were censored. Notice that the bounds get wider as the level of censoring increases. Moreover, censoring samples at the end of the event horizon increases the predictive uncertainty more than censoring samples in the beginning of it. Generally, fewer samples are available at the end of the event horizon, since most bearings have likely failed by then, so at this point, censoring is more detrimental to prediction accuracy.
@@
Figure 4: Predicted survival probability $S(t)$ using the KM estimator under various amounts of censoring. The shaded area around the curves represent empirical 95% confidence intervals, computed using the Greenwood formula [51]. We see that the confidence interval increases proportionally with the level of censoring, thus indicating more uncertainty in the survival probability.
Figure 5 shows a histogram of the censored and uncensored event times under various amounts of censoring. As the amount of censoring increases, the number of observed events becomes lower. Under 25% censoring, there are more observed events than censored events at each time point. Under 50% censoring, we see more censored samples than uncensored ones at later time points, i.e., between 0 and 20 minutes, since there are more samples with lower event times than higher. Under 75% censoring, this phenomenon is even more pronounced.
@@ 4.6 Survival models
The main assumption behind this model is the proportional hazards assumption, which states that the hazard ratios between instances remain constant over time. In other words, the effect of any given feature on the hazard function is multiplicative and does not vary with time.
@@
GBSA [50]: GBSA is a shallow survival model that adopts a gradient boosting strategy to estimate parameters $\bm{\theta}$ by maximizing Eq. 8 and considers a flexible set of candidate features. In each boosting step, there are $q_{k}$ predetermined candidate sets of features, and for each of these $q_{k}$ sets, it simultaneously updates the parameter for the corresponding feature. Regular gradient boosting would update only one component of $\bm{\theta}$ in component-wise boosting, or fit the gradient by using all features in each step. GBSA assumes proportional hazards.
@@
RSF [24]: The Random Survival Forests model is a shallow survival model extension of decision trees for survival analysis, which are made to handle censored data. During training, data are recursively partitioned based on some splitting criterion, and similar data points based on the event of interest are put on the same node. This ensures a decorrelation of individual trees by iteratively building each tree on a different bootstrap sample of the original training data and then at each node only evaluate the split criterion for a randomly selected subset of features and thresholds. Predictions are then formed by aggregating predictions of individual trees in the ensemble, like the regular Random Forests model. RSF does not assume proportional hazards, but rather that sampling is representative, i.e., by drawing a random number of samples, we would obtain an equal split between the number of outputs.
@@
MTLR [72]: Multi-Task Logistic Regression is a deep survival model that directly models the individual survival function by combining multiple local logistic regression models. It splits the discretized event horizon into $K$ bins $(\tau_{0}=0,\tau_{1}),(\tau_{1},\tau_{2}),\ldots,(\tau_{K-1},\tau_{K})$ and computes the probability of event at each bin by fitting a logistic regression model [45]. The predicted time to event $\hat{t_{i}}$ can be then be encoded deterministically using a $K$-dimensional binary sequence $\mathbf{y}_{i}=(y_{i,1},y_{i,2},\ldots,y_{i,K})\in\mathbb{Z}^{K}$, where $y_{i,k}=0$ means that the observation had not experienced the event at time $\tau_{k-1}$, and $y_{i,k}=1$ means that the observation experienced the event before time $\tau_{k-1}$ [45]. MTLR does not assume proportional hazards.
@@
BNNSurv [34]: BNNSurv is a deep survival model that trains a Bayesian Neural Network to predict risk scores. It uses the Cox partial log-likehood to optimize the objective Eq. (8), but the weights and biases of the network are treated as random variables with an unknown posterior distribution. This allows representing the uncertainty in the parameter distribution, as opposed to deterministic models, where the weights and biases are fixed numbers. During training, they are estimated with an approximate distribution using, e.g., variational inference (VI), Monte Carlo Dropout (MCD), or another inference technique. The output nodes thus provide the mean and standard deviation of a risk score, $r_{i}$, as Gaussian samples, that is, $r_{i}\sim\mathcal{N}\left(\mu_{\bm{x}_{i}},\sigma_{\bm{x}_{i}}\right)$ for a $d$-dimensional sample $\bm{x}_{i}$. In this work, we adopt the BNNSurv model using MCD inference and set a dropout rate of 25%. BNNSurv is based on CoxPH, and therefore assumes proportional hazards between the features and the risk scores.
@@ 5.1 Setup
As described, we implement the traditional Cox model (CoxPH) [10], GBSA [50], Random Survival Forests (RSF) [24], and two neural network-based models, Multi-Task Logistic Regression (MTLR) [72] and BNNSurv using Monte-Carlo Dropout (MCD) [34]. To estimate the generalization error of each model, we arrange a stratified 5-fold cross-validation procedure, ensuring that event times and censoring levels remain consistent across all folds. For each fold, we apply $z$-score data normalization to all features and compute the event horizon on the training set. We then report the average model performance across the five folds, evaluating both predictive performance and calibration. If the respective model requires a validation set for early stopping, we allocate 30% of the training set for this purpose. For the continuous-time models, i.e., CoxPH, GBSA, RSF, and BNNSurv, we compute the event horizon using the unique events and censoring times in the training set, thus the number of bins is the number of unique time points. For the discrete-time model, i.e., MTLR, we calculate the square root of the number of uncensored observations, $\sqrt{\sum_{i:\delta_{i=1}}\mathbbm{1}}$, in the training set and use quantiles to evenly divide these uncensored observations into bins, as in [21].
We adopt sensible hyperparameters for the survival models based on the specific operating condition, as each condition has a distinct dataset with a different number of samples. For the baseline CoxPH model, we set the convergence criterion to $10^{-9}$ and a maximum of 100 iterations for all condition variants. For the GBSA and RSF models, aligned with the three operating conditions, we set the number of estimators to $\{100,200,400\}$, the minimum number of samples needed to split to $\{5,10,20\}$, the minimum number of samples in a leaf node to $\{3,5,10\}$ and a maximum depth of $\{3,5,10\}$. For the MTLR model, we use batch sizes of $\{32,64,128\}$, hidden sizes of $\{16,32,64\}$, a fixed dropout rate of 0.25, a learning rate of $8\times 10^{-5}$ using the Adam optimizer [29], a penalty term of 0.01, and 5000 training epochs as suggested by [45]. For the BNNSurv model, we use batch and network sizes identical to MTLR, but with a learning rate of 0.001 using Adam and 100 epochs of training.
@@ 5.3 Event detector performance
Table 5 shows a comparison between the annotated event times by the proposed event detection algorithm and the actual failure times in minutes. The average differences between the end of the recording and the estimated time of failure are 42.6%, 48.7%, and 56.9% for the high, medium and low operating conditions, respectively. All events are detected before the end of the recording. These results support the use of the proposed event detector, as it provides a good estimate of the time when each bearing starts to fail (i.e., the onset of degradation), but before the end of life, where the bearing physically stops working. Ideally, a failure event should be detected well in advance of end of life for predictive maintenance to have an effect.
It is also worth noting that the bearings in the XJTU-SY dataset fail at entirely different times, even for the same operating condition. This points to other factors, e.g. temperature, lubrication, and previous wear having a significant influence on the lifetime. We do see an increasing average error for lighter loads, which can be further mitigated by adjusting $\lambda_{KL}$ for the event detector (see Algorithm 1), to ensure a proper threshold for the KL divergence. In an industrial setting, bearings are normally run under a lot less stress than in this dataset, so the event detector would have to be recalibrated on any new dataset, keeping in mind the operational overhead predictive maintenance has, i.e., what is the cost of not detecting a fault in a malfunctioning bearing (false negative), as opposed to detecting a fault in a healthy bearing (false positive).
@@ 5.4 Survival model performance
Table 6 shows the mean and standard deviation of the MAE, i.e., the predictive performance, averaged over five cross-validation folds. Generally, we see lower MAE under high load, where the event horizon is shorter, the bearings deteriorate faster, and the event times are much more consistent (see Table 5 and Figure 5), compared to medium and low load. Under high load, increasing the level of censoring does not affect MAEH, but leads to worse MAEM and MAEPO results, particularly at high levels of censoring ($C=75\%$). Both MAEM and MAEPO compute pseudo-observations [3] for censored observations, and more censoring increases the number of these synthetic observations. With more censoring comes more uncertainty, and consequently, the metric has to compute simulated event times for more samples, which generally leads to higher predictive errors. This trend is common for all models. The baseline CoxPH model gives good MAE scores under high load, but cannot keep up with its ensemble and neural network-based variants when the load decreases and the number of samples is larger, only exceeding the GBSA model. The RSF model overall has the lowest MAE across operating conditions and censoring levels, even outperforming the neural networks, and its MAEH score is stable across censoring levels. However, the main trend is that censoring generally decreases predictive accuracy, as censoring leads to a loss of information, which can make it more difficult to accurately estimate the survival or the hazard function. This appears predominantly in the MAEH and MAEPO results.
@@
Table 8 reports the true MAE for the respective datasets, models, and censoring levels. In this case, the true MAE is calculated by removing the censoring from the test set, i.e., decensoring it, and keeping the censoring in the training set. Concerning the results, RSF and BNNSurv give the lowest true MAE results on average across all datasets. We also see that increasing censoring leads to worse predictive accuracy across the board. This indicates that higher censoring levels make it more difficult for the models to accurately predict the correct outcome, leading to larger predictive errors. As a baseline comparison, we trained a linear model with an $\ell_{1}$ regularization term (LASSO) on the uncensored observations only, i.e.$\mathcal{D}^{\delta_{i=1}}$, thus disregarding the censored ones. Notably, we see that disregarding the censored data leads to poorer predictive accuracy, when comparing LASSO with the censoring-aware models, across the majority of datasets and censoring levels. This underlines the importance of incorporating censored observations into the model, as censored data still offer valuable information about the likelihood of an event occurring, even if the event has not yet taken place.
Table 9 shows the calibration results in terms of D-calibration and C-calibration. To measure D-calibration, we slice the predicted survival function, $S(t)$, into 10% intervals and accept the null hypothesis that the set $\{S_{i}(d_{i})\}$ is uniformly distributed between $[0,1]$ for all individual survival distributions $d_{i}$ if the $p$-value of a Pearson’s $\chi^{2}$ test is above 0.05. Thus, a calibrated survival model has $p$-values greater than 0.05. We can see from the table that all models pass the test in five out of five folds under high load, except BNNSurv, passing the test in four out of five folds. As we decrease the load and the event horizon becomes wider, more models fail the D-calibration test. MTLR is the only model that gives D-calibrated survival curves in all datasets, no matter the level of censoring. To measure C-calibration specifically for the BNNSurv model, we check if the predicted credible intervals match the observed probability intervals. The BNNSurv model is C-calibrated under 25% censoring in all five folds, but calibration performance decreases with higher censoring.

### v3

@@ 4.4 Data preprocessing
The XJTU-SY dataset is a tabular dataset, where the rows represent temporal feature values, and the columns represent feature names. There are five bearings per operating condition and three operating conditions in total (see Table 2). As a preliminary preprocessing step, we convert the temporal dataset to a supervised learning dataset by computing a rolling average with adjustable window length ($w$) and lag ($l$) based on the operating condition. For bearings under high load (C1), we use $w=2$ and $l=-1$. For medium load (C2), we use $w=4$ and $l=-3$. For the light load (C3), we use $w=6$ and $l=-5$. This reflects the length of the event horizon and means that more samples are gathered from bearings running for longer, but also that the averaging window is larger and the lag is higher. The moving window averages both the features and the event times. This produces three supervised learning datasets, $\mathcal{D}_{C1}$ ($N=635$, $d=12$), $\mathcal{D}_{C2}$ ($N=1586$, $d=12$) and $\mathcal{D}_{C3}$ ($N=4930$, $d=12$), where $N$ is the number of samples and $d$ is the number of features.
@@ 4.5 Data censoring
Censored observations, i.e., where the event of failure is not observed, are common in an industrial setting. Some bearings may not have failed yet at the time of observation, or they may have been decommissioned before failure occurred. Because the XJTU-SY dataset only includes complete observations, i.e., the failure time is always known, we introduce random and independent censoring by a fixed percentage, $C$, to simulate censoring in an industrial setting. The amount of censoring in an industrial setting can vary and depends on several factors, e.g., some bearings may be decommissioned in advance to avoid failure or perform preventive maintenance. To evaluate our method under various realistic amounts of censoring, we selected 25%, 50%, and 75%, resulting in three censored data sets with the respective level of censoring. For each of these datasets, we randomly sampled the defined percentages of samples from the dataset $\mathcal{D}$, and designated those as censored. For example, to censor an observation, $i$, we calculate the time of censoring, $c_{i}$, as a random number between one and the observed event time, $e_{i}$, and use $c_{i}$ instead of $e_{i}$ as the time-to-event for observation $i$, i.e.$t_{i}=c_{i}$ and $\delta_{i}=0$. This indicates that the bearing was censored at some time between when it was recorded and when it actually failed. See B for a pseudocode implementation of the algorithm that automatically performs these changes.
Figure 4 shows the KM estimate of the survival function after random censoring was applied to the individual datasets. The shaded area around the curve represents an empirical 95% confidence interval with upper and lower bounds. These bounds are the best/worst cases of the survival function given the observed data without any parametric assumptions. We calculate the bounds using the Greenwood formula [28], and they represent the predictive uncertainty and thus the consequences of censoring when making estimates. For the upper bound on the survival function, we assume that all censored bearings do not experience the event. For the lower bound, we assume that all censored bearings immediately experience the event after they were censored. Notice that the bounds get wider as the level of censoring increases. Moreover, censoring samples at the end of the event horizon increases the predictive uncertainty more than censoring samples at the beginning of it. Generally, fewer samples are available at the end of the event horizon since most bearings have probably failed by then, so at this point, censoring is more detrimental to prediction accuracy.
@@
Figure 4: Predicted survival probability $S(t)$ using the KM estimator under various amounts of censoring. The shaded area around the curves represent empirical 95% confidence intervals, computed using the Greenwood formula [sawyer2003greenwood]. We see that the confidence interval increases proportionally with the level of censoring, thus indicating more uncertainty in the survival probability.
Figure 5 shows a histogram of censored and uncensored event times under various amounts of censoring. As censoring increases, the number of observed events decreases. Under 25% censoring, there are more observed events than censored events at each time point. Under 50% censoring, we see more censored samples than uncensored ones at later time points, i.e., between 0 and 20 minutes, since there are more samples with lower event times than higher. Under 75% censoring, this phenomenon is even more pronounced.
@@ 4.6 Survival models
The main assumption behind this model is the proportional hazards assumption, which states that the hazard ratios between instances remain constant over time. In other words, the effect of any given feature on the hazard function is multiplicative and does not vary over time.
@@
GBSA [52]: GBSA is a shallow survival model that adopts a gradient boosting strategy to estimate parameters $\bm{\theta}$ by maximizing Eq. 8 and considers a flexible set of candidate features. In each boosting step, there are $q_{k}$ predetermined candidate sets of features, and for each of these $q_{k}$ sets, it simultaneously updates the parameter for the corresponding feature. Regular gradient boosting would update only one component of $\bm{\theta}$ in component-wise boosting, or fit the gradient using all features in each step. GBSA assumes proportional hazards.
@@
RSF [25]: The Random Survival Forests model is a shallow survival model extension of decision trees for survival analysis, which are made to handle censored data. During training, data are recursively partitioned according to some splitting criterion, and similar data points based on the event of interest are placed on the same node. This ensures a decorrelation of individual trees by iteratively building each tree on a different bootstrap sample of the original training data and then at each node only evaluate the split criterion for a randomly selected subset of features and thresholds. Predictions are then made by aggregating the predictions of the individual trees in the ensemble, like the regular Random Forests model [7]. RSF does not assume proportional hazards, but rather that sampling is representative, i.e., by drawing a random number of samples, we would obtain an equal split between the number of outputs.
@@
MTLR [73]: Multi-Task Logistic Regression is a deep survival model that directly models the individual survival function by combining multiple local logistic regression models. It splits the discretized event horizon into $K$ bins $(\tau_{0}=0,\tau_{1}),(\tau_{1},\tau_{2}),\ldots,(\tau_{K-1},\tau_{K})$ and computes the probability of event at each bin by fitting a logistic regression model [47]. The predicted time-to-event $\hat{t_{i}}$ can be then be encoded deterministically using a $K$-dimensional binary sequence $\mathbf{y}_{i}=(y_{i,1},y_{i,2},\ldots,y_{i,K})\in\mathbb{Z}^{K}$, where $y_{i,k}=0$ means that the observation had not experienced the event at time $\tau_{k-1}$, and $y_{i,k}=1$ means that the observation experienced the event before time $\tau_{k-1}$ [47]. MTLR does not assume proportional hazards.
@@
BNNSurv [36]: BNNSurv is a deep survival model that trains a Bayesian Neural Network to predict risk scores. It uses the Cox partial log-likehood to optimize the objective Eq. (8), but the weights and biases of the network are treated as random variables with an unknown posterior distribution. This allows representing the uncertainty in the parameter distribution, as opposed to deterministic models, where the weights and biases are fixed numbers. During training, they are estimated with an approximate distribution using e.g., variational inference (VI), Monte Carlo dropout (MCD) or another inference technique. The output nodes thus provide the mean and standard deviation of a risk score, $r_{i}$, as Gaussian samples, that is, $r_{i}\sim\mathcal{N}\left(\mu_{\bm{x}_{i}},\sigma_{\bm{x}_{i}}\right)$ for a $d$-dimensional sample $\bm{x}_{i}$. In this work, we adopt the BNNSurv model using MCD inference and set a dropout rate of 25%. BNNSurv is based on CoxPH and, therefore, assumes proportional hazards between the features and the risk scores.
@@ 5.1 Setup
As described, we implement the traditional Cox model (CoxPH) [11], GBSA [52], Random Survival Forests (RSF) [25], and two neural network-based models, Multi-Task Logistic Regression (MTLR) [73] and BNNSurv using Monte-Carlo Dropout (MCD) [36]. To estimate the generalization error of each model, we arrange a stratified 5-fold cross-validation procedure, ensuring that event times and censoring levels remain consistent across all folds. For each fold, we apply $z$-score data normalization to all features and compute the event horizon on the training set. We then report the average model performance across the five folds, evaluating both predictive performance and calibration. If the respective model requires a validation set for early stopping, we allocate 30% of the training set for this purpose. For continuous-time models, i.e., CoxPH, GBSA, RSF, and BNNSurv, we compute the event horizon using the unique events and censoring times in the training set, thus the number of bins is the number of unique time points. For the discrete-time model, i.e., MTLR, we calculate the square root of the number of uncensored observations, $\sqrt{\sum_{i:\delta_{i=1}}\mathbbm{1}}$, in the training set and use quantiles to evenly divide these uncensored observations into bins, as in [22].
We adopt sensible hyperparameters for the survival models based on the specific operating condition, as each condition has a distinct dataset with a different number of samples. For the baseline CoxPH model, we set the convergence criterion to $10^{-9}$ and a maximum of 100 iterations for all conditions. For the GBSA and RSF models, aligned with the three operating conditions, we set the number of estimators to $\{100,200,400\}$, the minimum number of samples needed to split to $\{5,10,20\}$, the minimum number of samples in a leaf node to $\{3,5,10\}$ and a maximum depth of $\{3,5,10\}$. For the MTLR model, we use batch sizes of $\{32,64,128\}$, hidden sizes of $\{16,32,64\}$, a fixed dropout rate of 0.25, a learning rate of $8\times 10^{-5}$ using the Adam optimizer [31], a penalty term of 0.01, and 5000 training epochs as suggested by [47]. For the BNNSurv model, we use batch and network sizes identical to MTLR, but with a learning rate of 0.001 using Adam and 100 epochs of training.
@@ 5.3 Event detector performance
Table 5 shows a comparison between the annotated event times by the proposed event detection algorithm and the actual failure times in minutes. The average differences between the end of the recording and the estimated time of failure are 42.6%, 48.7%, and 56.9% for the high, medium and low operating conditions, respectively. All events are detected before the end of the recording. These results support the use of the proposed event detector, as it provides a good estimate of the time when each bearing starts to fail (i.e., the onset of degradation), but before the end of life, when the bearing physically stops working. Ideally, a failure event should be detected well in advance of the end of life for predictive maintenance to have an effect.
It is also worth noting that the bearings in the XJTU-SY dataset fail at entirely different times, even for the same operating conditions. This points to other factors – e.g., temperature, lubrication, and previous wear having a significant influence on the lifetime. We do see an increasing average error for lighter loads, which can be further mitigated by adjusting $\lambda_{KL}$ for the event detector (see Algorithm 1), to ensure a proper threshold for the KL divergence. In an industrial setting, bearings are normally run under less stress than in this dataset, so the event detector would have to be recalibrated in any new dataset, keeping in mind the operational overhead predictive maintenance has, i.e., what is the cost of not detecting a fault in a malfunctioning bearing (false negative), as opposed to detecting a fault in a healthy bearing (false positive).
@@ 5.4 Survival model performance
Table 6 shows the mean and standard deviation of the MAE, i.e., the predictive performance, averaged over five cross-validation folds. Generally, we see lower MAE under high load, where the event horizon is shorter, the bearings deteriorate faster, and the event times are much more consistent (see Table 5 and Figure 5), compared to medium and low load. Under high load, increasing the level of censoring does not affect MAEH, but leads to worse MAEM and MAEPO results, particularly at high levels of censoring ($C=75\%$). Both MAEM and MAEPO compute pseudo-observations [3] for censored observations, and more censoring increases the number of these synthetic observations. With more censoring comes more uncertainty, and consequently, the metric has to compute simulated event times for more samples, which generally leads to higher predictive errors. This trend is common for all models. The baseline CoxPH model gives good MAE scores under high load, but cannot keep up with its ensemble and neural network-based variants when the load decreases and the number of samples is larger, only exceeding the GBSA model. In general, the RSF model has the lowest MAE in operating conditions and in censoring levels, even outperforming neural networks, and its MAEH score is stable in all levels of censoring. However, the main trend is that censoring generally decreases predictive accuracy, as censoring leads to a loss of information, which can make it more difficult to accurately estimate the survival or the hazard function. This appears predominantly in the MAEH and MAEPO results.
@@
Table 8 reports the true MAE for the respective datasets, models, and censoring levels. In this case, the true MAE is calculated by removing the censoring from the test set, i.e., decensoring it, and keeping the censoring in the training set. Concerning the results, RSF and BNNSurv give the lowest true MAE results on average across all datasets. We also see that increasing censoring leads to worse predictive accuracy across the board. This indicates that higher levels of censoring make it more difficult for models to accurately predict the correct result, leading to larger errors. As a baseline comparison, we trained a linear model with an $\ell_{1}$ regularization term (LASSO) on uncensored observations only, i.e.$\mathcal{D}^{\delta_{i=1}}$, thus disregarding censored ones. In particular, we see that disregarding the censored data leads to poorer predictive accuracy (i.e., MAE), when comparing LASSO with the censoring-aware models, across the majority of datasets and censoring levels. This underlines the importance of incorporating censored observations into the model, as censored data still offer valuable information about the likelihood of an event occurring, even if the event has not yet taken place.
Table 9 shows the calibration results in terms of D-calibration and C-calibration. To measure D-calibration, we slice the predicted survival function, $S(t)$, into 10% intervals and accept the null hypothesis that the set $\{S_{i}(d_{i})\}$ is uniformly distributed between $[0,1]$ for all individual survival distributions $d_{i}$ if the $p$ value of a Pearson test $\chi^{2}$ is greater than 0.05. Thus, a calibrated survival model has $p$-values greater than 0.05. We can see from the table that all models pass the test in five out of five folds under high load, except BNNSurv, which passes the test in four out of five folds. As we decrease the load and the event horizon becomes wider, more models fail the D-calibration test. MTLR is the only model that gives D-calibrated survival curves in all datasets, regardless of the level of censoring. To measure C-calibration specifically for the BNNSurv model, we check if the predicted credible intervals match the observed probability intervals. The BNNSurv model is C-calibrated under 25% censoring in all five folds, but the calibration performance decreases with higher censoring.